### v1

# Efficient Algorithms for Computing the Sensitivity of the Pearson Correlation Coefficient to New Data

Abstract
Reliably measuring the collinearity of bivariate data is crucial in statistics, particularly for time-series analysis or ongoing studies in which incoming observations can significantly impact current collinearity estimates. To this end, we introduce a novel algorithm for computing the sensitivity of the Pearson correlation coefficient (PCC) to new and potentially outlying data points, subject to straightforward constraints. Leveraging identities from Welford’s online algorithm for sample variance, we develop a rigorous theoretical framework for analyzing the maximal change to the PCC that can be induced by additional data. Further, we show that the resulting optimization problem yields elegant closed-form solutions that can be accurately computed by linear- and constant-time algorithms. Our work not only creates new theoretical avenues for robust correlation measures, but also has broad practical implications for disciplines that span econometrics, operations research, clinical trials, climatology, differential privacy, and bioinformatics. Software implementations of our algorithms are made available on GitHub both for practical deployment and future theoretical development.
1 Introduction
The Pearson Correlation Coefficient (PCC), one of the simplest measures of collinearity between two random variables, has been an essential component of data analysis for over a century [23]. Along with other collinearity measures, it finds applications in quantitative disciplines that span physics, psychology, economics, and mechanical engineering [1]. Like many statistical quantities, however, its utility can be significantly curtailed by the presence of outliers—data points that differ widely from the underlying distribution from which the majority of a dataset is sampled [14]. This issue is exacerbated by online data collection, during which new—and potentially outlying—data points have the potential to significantly alter the overall sample PCC.
Such scenarios are common across fields that rely on time-series data or otherwise ongoing observations, necessitating an understanding of the extent to which online PCC estimates are susceptible to change. For example, in medicine, linear relationships between patient covariates and clinical outcomes are often continuously monitored during drug development trials [27, 7, 15]. In finance and econometrics, investors dynamically manage portfolio risk via tools that include the PCC, but are also forced to adjust these estimates given changing market conditions, evolving relationships between financial instruments, and ongoing data collection [10, 12, 17].
Obviously, frequentist statistics provides many common methods to quantify the certainty of sample estimates, including standard confidence intervals and hypothesis-testing [1]. As for the problem of accounting for outliers, robust statistics [14, 19] offers both simple and sophisticated analytical methods for estimating parameters that might be affected by outlying data. For instance, winsorizing, analogous to clipping in signal processing, offers a straightforward technique for minimizing the effect of outliers [6]. Order statistics [24], and their generalizations in the form of M-statistics and other measures [14, 19], offer a high breakdown-point. This refers to the fraction of data points that are required to be outliers before the metrics’ values extend beyond the range of inlying data [14, 19]. For the specific case of linear relationships, [26] developed efficient trimming methods like Least Trimmed Squares (LTS) for rejecting outliers in regression analysis. Beyond the breakdown point, quantifying the affect of outliers on data has also been studied extensively via measures like Cook’s distance [2] and empirical impact functions [9].
However, none of these approaches either considers the worst-case change in collinearity measures as a result of online collection or offers an elegant method for computing it. Measures of statistical significance like p-values also evolve as a function of incoming data, meaning that if the desired metric is not the PCC itself but rather its associated p-value, hypothesis-testing does not address the challenge of learning the target metric online. While robust statistics can provide invaluable insight into the effect of outliers on datasets, it assumes that all data points are fixed and that outliers are already present in the sample [14, 19]. Moreover, many robust estimators can be quite intensive to compute [14, 19], and unlike their far more popular and common non-robust counterparts like the PCC, are less immediately interpretable. As is the case for robust statistics, influence measures like empirical impact functions are also intended for scenarios in which outliers already exist, rather than for dynamic settings in which statisticians must simultaneously consider the set of all potential values that can subsequent observations can take.
In the following work, we derive analytical solutions to the sensitivity of the PCC of datasets with respect to the addition of new points, including for the worst-case observations that will induce the maximal change. By drawing from Welford’s online algorithm for computing variance [31], we derive a rigorous framework for analyzing the sensitivity of online collinearity measures. Our main results are that the solutions for the worst-case points are not only closed-form, but easily computable in either linear or constant time with a high degree of numerical precision.
This paper is structured as follows. In Section 2, we establish prerequisite definitions, clarify notation, and state preliminary theorems. The primary results of this work are in Section 3, in which we develop the closed-form solutions for primary sensitivity. Section 4 provides pseudocode, a proof of correctness, and runtime analysis for an algorithm to efficiently compute these solutions. Experiments are conducted on synthetic and real-world data in Section 5, which are then followed by a brief discussion and suggestions for future directions in Section 6. We conclude in Section 7.
2 Prerequisites
Table 1: Key notation
Symbol Description
Topology of Feasible Region
$\mathcal{F}$ Feasible region
$\text{int}\left(\mathcal{F}\right)$ Interior of feasible region
$\partial\mathcal{F}$ Boundary of feasible region
$\partial\mathcal{U}_{y},\partial\mathcal{L}_{y},\partial\mathcal{U}_{x},\partial\mathcal{L}_{x}$ Upper, lower, right, left edges of $\partial\mathcal{F}$
$\ell_{x},u_{x},\ell_{y},u_{y}$ Lower, upper bounds for $x$, $y$
$\mathcal{C}=\left\{C_{\ell,\ell},C_{\ell,u},C_{u,\ell},C_{u,u}\right\}$ Corner points
$\mathcal{I}_{n-1}=\left\{I_{x,\ell,n-1},I_{x,u,n-1},I_{y,\ell,n-1},I_{y,u,n-1}\right\}$ Intersection points
Welford’s Quantities
$s_{x,n-1},s_{y,n-1}$ Biased sample standard deviations of $D$
$s_{xy,n-1}$ Biased sample covariance of $D$
$\bar{x}_{n-1},\bar{y}_{n-1}$ Sample means of $D$
Statistical Measures
$D$ Bounded multiset
$n-1$ Number of points in $D$
$\left(x_{i},y_{i}\right)$ Individual point in $D$
$\left(x_{n},y_{n}\right)$ New data point to add to $D$
$r_{n-1}\equiv\textnormal{cor}(D)$ PCC of $D$
$r_{n}\left(x_{n},y_{n}\right)\equiv\textnormal{cor}\left(D\cup\left\{(x_{n},y_{n})\right\}\right)$ PCC of $D$ with new point
$S_{1}(D,\mathcal{F})$ Primary sensitivity of $D$ within $\mathcal{F}$
$S_{k}(D,\mathcal{F})$ $k$-ary sensitivity of $D$ within $\mathcal{F}$
Regression Parameters
$\hat{\alpha}_{x,n-1},\hat{\beta}_{x,n-1},\hat{\alpha}_{y,n-1},\hat{\beta}_{y,n-1}$ BLUE parameters regressing $y$ onto $x$, $x$ onto $y$
$\mathcal{G}_{x,n-1},\mathcal{G}_{y,n-1}$ Sets of points lying on BLUE lines
Miscellaneous
$\mathbf{H}_{r_{n}}\left(x_{n},y_{n}\right)$ Hessian matrix of $r_{n}$ at $\left(x_{n},y_{n}\right)$
$\cup$ Union operator (for both sets and multisets)
||$\cdot$|| ⋅ | Cardinality (for both sets and multisets)
Concretely, we consider a bounded dataset
$D=\{(x_{i},y_{i})\}_{i=1}^{n-1}\quad\text{where}\quad\ell_{x}\leq x_{i}\leq u_{x}\quad\text{and}\quad\ell_{y}\leq y_{i}\leq u_{y}\quad\forall i$ (1)
and put the feasible region (see Figure 1)
$\mathcal{F}:=\left[\ell_{x},u_{x}\right]\times\left[\ell_{y},u_{y}\right].$ (2)
Note that in this section, for brevity, it is easier to define $n$ such that $|D|=n-1$. We also note that $D$ is not a set in the traditional sense, but rather a multiset; moreover, we abuse notation slightly by indexing each point $(x_{i},y_{i})$ and by using superscripts to indicate the Cartesian product of $\mathcal{F}$ with itself in the context of multisets. We define the sample correlation coefficient of $D$ as follows.
Definition 2.1.
The Pearson correlation coefficient (PCC) of a dataset $D$ is given by
$\textnormal{cor}\left(D\right):=\frac{\sum_{i=1}^{n-1}\left(x_{i}-\bar{x}_{n-1}\right)\left(y_{i}-\bar{y}_{n-1}\right)}{\sqrt{\sum_{i=1}^{n-1}\left(x_{i}-\bar{x}_{n-1}\right)^{2}\sum_{i=1}^{n-1}\left(y_{i}-\bar{y}_{n-1}\right)^{2}}},$ (3)
where $\bar{x}_{n-1}$ and $\bar{y}_{n-1}$ are the sample means of $D$ [23].
This definition makes our motivation for imposing constraints on our datasets more obvious. Clearly, if incoming observations were not bounded, worst-case points could be arbitrarily large, overshadowing the present data and skewing the PCC to an arbitrary degree towards -1 or 1. We also opt for a rectangular feasible region due its simplicity and similarity to winsorization. However, we consider more complex boundaries below in Section 6.
In online settings, we consider the effect of adding a new point $\left(x_{n},y_{n}\right)$ to $D$ on the PCC, with an interest in the robustness of the PCC to the worst case. We formalize this as follows.
Definition 2.2.
We call
$S_{1}\left(D,\mathcal{F}\right):=\max_{\left(x_{n},y_{n}\right)\in\mathcal{F}}\lvert\textnormal{cor}\left(D\right)-\textnormal{cor}\left(D\cup\left\{\left(x_{n},y_{n}\right)\right\}\right)\rvert$ (4)
the primary sensitivity of $D$ within $\mathcal{F}$.
For the more general case of adding up to $k$ points to $D$, we define the $k$-ary sensitivity, which will return to briefly in Section 6.
Definition 2.3.
We call
$S_{k}\left(D,\mathcal{F}\right):=\max_{D^{\prime}\in\mathcal{F}^{k}}\lvert\textnormal{cor}\left(D\right)-\textnormal{cor}\left(D\cup D^{\prime}\right)\rvert$ (5)
the $k$-ary sensitivity of $D$ within $\mathcal{F}$.
Drawing from Welford’s online algorithm for computing variance and covariance, we leverage the following recurrence relations between $D$ and subsequent data points [31].
Theorem 2.4 (Welford’s recurrence relations).
Fix a dataset $D=\{(x_{i},y_{i})\}_{i=1}^{n-1}$ and new point $\left(x_{n},y_{n}\right)$. Let $s_{x,n-1}$ and $\bar{x}_{n-1}$ be the biased sample standard deviations and sample means, respectively, of $\left\{x_{1},x_{2},\ldots,x_{n-1}\right\}$. Further, let $s_{x,n}$ and $\bar{x}_{n}$ be the sample mean of $\left\{x_{1},x_{2},\ldots,x_{n-1},x_{n}\right\}$. Then
$\bar{x}_{n}=\frac{(n-1)\bar{x}_{n-1}+x_{n}}{n}$ (6)
and
$s^{2}_{x,n}=\frac{(n-1)s^{2}_{x,n-1}+(x_{n+1}-\bar{x}_{n-1})(x_{n+1}-\bar{x}_{n})}{n},$ (7)
Moreover,
$s_{xy,n}=\frac{n-1}{n}\left(s_{xy,n-1}+\frac{(x_{n}-\bar{x}_{n-1})(y_{n}-\bar{y}_{n-1})}{n}\right),$ (8)
where $s_{xy,n}$ denotes sample covariance [31].
We will also find that the best linear unbiased estimators (BLUE) for the regression line for a data set $D$ play a role in proving our key theorems [1]. The BLUE parameters can be related to the statistical moments computed by Welford via the Gauss-Markov Theorem [1].
Theorem 2.5 (Gauss-Markov Theorem).
Fix a dataset $D=\{(x_{i},y_{i})\}_{i=1}^{n-1}$. Given a model function
$y_{i}=\alpha_{x,n-1}+\beta_{x,n-1}x_{i}$ (9)
regressing $\left\{y_{1},y_{2},\ldots,y_{n-1}\right\}$ onto $\left\{x_{1},x_{2},\ldots,x_{n-1}\right\}$, the least-squares line is given by
$\hat{\beta}_{x,n-1}=\frac{s_{xy,n-1}}{s_{x,n-1}^{2}}\quad\text{ and }\quad\hat{\alpha}_{x,n-1}=\bar{y}_{n-1}-\beta_{x,n-1}\bar{x}_{n-1}.$ (10)
Remark.
As is often noted in proofs of the Gauss-Markov Theorem, we observe that the BLUE lines intersect at the point $(\bar{x}_{n-1},\bar{y}_{n-1})$ and that the sets of points $\mathcal{G}_{x,n-1}$ and $\mathcal{G}_{y,n-1}$ are equal if and only if $\textnormal{cor}\left(D\right)=\pm 1$.
Finally, the following points will play a critical role in our solutions to primary sensitivity.
Definition 2.6.
We refer to the points $\left(\ell_{x},\ell_{y}\right)$, $\left(\ell_{x},u_{y}\right)$, $\left(u_{x},\ell_{y}\right)$, and $\left(u_{x},u_{y}\right)$, which we denote $C_{\ell,\ell}$, $C_{\ell,u}$, $C_{u,\ell}$, and $C_{u,u}$, respectively, as corner points.
Definition 2.7.
We refer to the points $\left(\hat{\beta}_{x,n}^{-1}\left(\ell_{y}-\hat{\alpha}_{x,n}\right),\ell_{y}\right)$, $\left(\hat{\beta}_{x,n}^{-1}\left(u_{y}-\hat{\alpha}_{x,n}\right),u_{y}\right)$, $\left(\ell_{x},\hat{\beta}_{y,n}^{-1}\left(\ell_{y}-\hat{\alpha}_{y,n}\right)\right)$, and $\left(u_{x},\hat{\beta}_{y,n}^{-1}\left(u_{x}-\hat{\alpha}_{y,n}\right)\right)$, which we denote $I_{x,\ell,n-1}$, $I_{x,u,n-1}$, $I_{y,\ell,n-1}$, and $I_{y,u,n-1}$, respectively, as intersection points.
Figure 1: An illustration of the geometry of $\mathcal{F}$, $D$, and the intersection and corner points. In this example, only $I_{x,u,n-1}\in\mathcal{F}$.
3 Primary sensitivity
In the following section, we leverage Welford [1] to derive a closed-form function mapping from new points $\left(x_{n},y_{n}\right)$ to the quantity $\textnormal{cor}\left(D\cup\left\{\left(x_{n},y_{n}\right)\right\}\right)$. By studying the solutions to this function, we will then be able to trivially compute $S_{1}$. Substituting the definition of correlation, and rewriting it as a function $r_{n}:\mathcal{F}\to\left[-1,1\right]$, we have
$r_{n}(x_{n},y_{n})$ $=\frac{s_{xy,n-1}}{s_{x,n-1}s_{y,n-1}}$
$=\frac{nr_{n-1}s_{x,n-1}s_{y,n-1}+\left(x_{n}-\bar{x}_{n-1}\right)\left(y_{n}-\bar{y}_{n-1}\right)}{\sqrt{\left(ns_{x,n-1}^{2}+\left(x_{n}-\bar{x}_{n-1}\right)^{2}\right)\left(ns_{y,n-1}^{2}+\left(y_{n}-\bar{y}_{n-1}\right)^{2}\right)}}.$ (11)
For brevity, we let $s_{x}=s_{x,n-1}$, $s_{y}=s_{y,n-1}$, $\Delta x=x_{n}-\bar{x}_{n-1}$, $\Delta y=y_{n}-\bar{y}_{n-1}$, $r=r_{n-1}$, then observe that we have the following derivatives of $r_{n}$:
$\frac{\partial r_{n}}{\partial x_{n}}$ $=\frac{ns_{x}\left(s_{x}\Delta y-rs_{y}\Delta x\right)}{\left(ns_{x}^{2}+\Delta x^{2}\right)^{\frac{3}{2}}\left(ns_{y}^{2}+\Delta y^{2}\right)^{\frac{1}{2}}}$ (12)
$\frac{\partial r_{n}}{\partial y_{n}}$ $=\frac{ns_{y}\left(s_{y}\Delta x-rs_{x}\Delta y\right)}{\left(ns_{x}^{2}+\Delta x^{2}\right)^{\frac{1}{2}}\left(ns_{y}^{2}+\Delta y^{2}\right)^{\frac{3}{2}}}$ (13)
$\frac{\partial^{2}r_{n}}{\partial x_{n}^{2}}$ $=\frac{ns_{x}\left(2rs_{y}\Delta x^{2}-nrs_{x}^{2}s_{y}-3s_{x}\Delta x\Delta y\right)}{\left(ns_{x}^{2}+\Delta x^{2}\right)^{\frac{5}{2}}\left(ns_{y}^{2}+\Delta y^{2}\right)^{\frac{1}{2}}}$ (14)
$\frac{\partial^{2}r_{n}}{\partial y_{n}^{2}}$ $=\frac{ns_{y}\left(2rs_{x}\Delta y^{2}-nrs_{x}s_{y}^{2}-3s_{y}\Delta x\Delta y\right)}{\left(ns_{x}^{2}+\Delta x^{2}\right)^{\frac{1}{2}}\left(ns_{y}^{2}+\Delta y^{2}\right)^{\frac{5}{2}}}$ (15)
$\frac{\partial^{2}r_{n}}{\partial x_{n}\partial y_{n}}$ $=\frac{ns_{x}s_{y}\left(ns_{x}s_{y}+r\Delta x\Delta y\right)}{\left(ns_{x}^{2}+\Delta x^{2}\right)^{\frac{3}{2}}\left(ns_{y}^{2}+\Delta y^{2}\right)^{\frac{3}{2}}}$ (16)
Proposition 3.1.
For all $\left(x_{n},y_{n}\right)\in\mathcal{F}$, $\frac{\partial r_{n}}{\partial x_{n}}=0$ $\left(\frac{\partial r_{n}}{\partial y_{n}}=0\right)$ if and only if $\left(x_{n},y_{n}\right)\in\mathcal{G}_{x,n-1}$ $\left(\left(x_{n},y_{n}\right)\in\mathcal{G}_{y,n-1}\right)$.
Proof.
It suffices to to prove our proposition for $x_{n}$. If $\frac{\partial r_{n}}{\partial x_{n}}=0$, we must have
$s_{x}\left(y_{n}-\bar{y}_{n-1}\right)-r_{n-1}s_{y}\left(x_{n}-\bar{x}_{n-1}\right)=0,$ (17)
meaning
$y_{n}$ $=r_{n-1}\frac{s_{y}}{s_{x}}\left(x_{n}-\bar{x}_{n-1}\right)+\bar{y}_{n-1}$
[…]
$=\hat{\beta}_{x,n-1}x_{n}-\hat{\beta}_{x,n-1}\bar{x}_{n-1}+\bar{y}_{n-1}$
$=\hat{\beta}_{x,n-1}x_{n}+\hat{\alpha}_{x,n-1}.$
Therefore, $\left(x_{n},y_{n}\right)\in\mathcal{G}_{x,n-1}$. Our proof of the converse proceeds identically. ∎
Proposition 3.2.
$\operatorname*{argmax}_{\mathcal{F}}r_{n},\ \operatorname*{argmin}_{\mathcal{F}}r_{n}\in\partial\mathcal{F}$ (18)
Proof.
Without loss of generality, suppose that we have a unique $(x_{n},y_{n})=\operatorname*{argmax}r_{n}$, and, by way of contradiction, that $\left(x_{n},y_{n}\right)\in\text{int}\left(\mathcal{F}\right)$. Because $r_{n}$ is continuous, it must take maximal and minimal values along the closed and bounded set $\mathcal{F}$. We then know that we must have
$\frac{\partial r_{n}}{\partial x_{n}}=\frac{\partial r_{n}}{\partial y_{n}}=0.$ (19)
Furthermore, we then know that $\left(x_{n},y_{n}\right)$ must lie at the intersection of both $\mathcal{G}_{x,n-1}$ and $\mathcal{G}_{y,n-1}$. We then have two cases:
  1. (i)
    $\mathcal{G}_{x,n-1}=\mathcal{G}_{y,n-1}$, or
  2. (ii)
    $\mathcal{G}_{x,n-1}\cap\mathcal{G}_{y,n-1}=\left\{\left(x_{n},y_{n}\right)\right\}$.
Suppose that (i). Then we have $\nabla r_{n}\left(x_{n},y_{n}\right)=(0,0)$ for all $\left(x_{n},y_{n}\right)\in\mathcal{G}_{x,n-1}=\mathcal{G}_{y,n-1}$, in which case $\mathcal{G}_{x,n-1}=\mathcal{G}_{y,n-1}$ must constitute a “ridge” along which the function $r_{n}$ is constant. Additionally, we observe that $\mathcal{G}_{x,n-1}=\mathcal{G}_{y,n-1}$ must intersect the boundary at some point $\left(x_{n}^{\prime},y_{n}^{\prime}\right)\in\partial\mathcal{F}$ because both sets are continuous. Then $r_{n}(x_{n}^{\prime},y_{n}^{\prime})=r_{n}(x_{n},y_{n})$, meaning that $(x^{\prime}_{n},y^{\prime}_{n})=\operatorname*{argmax}r_{n}$, which is a contradiction.
We must instead have (ii), meaning $\left(x_{n},y_{n}\right)=\left(\bar{x}_{n-1},\bar{y}_{n-1}\right)$. Therefore, to determine whether $\left(x_{n},y_{n}\right)$ is a local maximum, we perform the second partial derivative test. Evaluating the determinant of the Hessian, we find that
$\text{det}\left(\mathbf{H}_{r_{n}}\left(x_{n},y_{n}\right)\right)=\frac{r_{n-1}^{2}-1}{n^{2}s_{x}^{2}s_{y}^{2}}.$ (20)
Because we have (ii), we know that $r\in\left(-1,1\right)$, meaning that $\text{det}\left(\mathbf{H}_{r_{n}}\left(x_{n},y_{n}\right)\right)<0$. Then $\left(x_{n},y_{n}\right)$ is not a local maximum, which is a contradiction.
Thus, we must have $\operatorname*{argmax}r_{n}\in\mathcal{F}$, which completes our proof. ∎
Lemma 3.3.
$\operatorname*{argmax}_{\mathcal{F}}r_{n},\ \operatorname*{argmin}_{\mathcal{F}}r_{n}\in\mathcal{C}\cup\mathcal{I}_{n}.$ (21)
Proof.
We extend Proposition 3.2 by dividing $\mathcal{F}$ into its four edges. Without loss of generality, we consider the edge $\partial\mathcal{U}_{y}$ and extremum $\left(x_{n},y_{n}\right)$ and show that $\left(x_{n},y_{n}\right)\in\left\{C_{\ell,u},C_{u,u},I_{x,u,n-1}\right\}$. Fixing $x_{n}=u_{x}$, we treat $r_{n}$ as a one-dimensional function of $y_{n}$ on the closed interval $\left[\ell_{y},u_{y}\right]$. We know that $\left(x_{n},y_{n}\right)$ must lie on (i) the one-dimensional endpoint $\ell_{y}$ or $u_{y}$ or (ii) the stationary point at which $\frac{\partial r_{n}}{\partial y_{n}}=0$. If (i), then $\left(x_{n},y_{n}\right)=C_{\ell,u}$ or $\left(x_{n},y_{n}\right)=C_{u,u}$. If (ii), then by Proposition 3.1, we must have $\left(x_{n},y_{n}\right)=I_{x,u,n-1}$. This completes our proof. ∎
Figure 2: (a) The same dataset in Figure 1 with both BLUE lines superimposed on the function $r_{n}$. (b) A 3D view of $r_{n}$ for the same dataset in Figure 1. Critically, $\left(\bar{x}_{n-1},\bar{y}_{n-1}\right)$ is a saddle point. (c) A dataset similar to that in Figure 1, albeit such that $\textnormal{cor}\left(D\right)=1$. Both BLUE lines coincide. (d) A 3D view of the same dataset in Figure 1(c). The BLUE lines form a “ridge” along which $\nabla r_{n}=(0,0)$.
4 Main algorithm
We are now able to develop our key algorithm, ComputePrimarySensitivity. Note that for its implementation, we assume that $|D|=n$, not $n-1$ as we did above.
Theorem 4.1.
ComputePrimarySensitivity correctly returns $S_{1}$.
Proof.
This follows trivially from Lemma 3.3. ∎
Theorem 4.2.
ComputePrimarySensitivity runs in $O(n)$ time.
Proof.
Each of lines 2, 4-10, and 12-15 will require $O(1)$ time. The loop in line 11 will run for a constant number of iterations given that the number of corner and intersection points is bounded. Therefore, the algorithm is dominated by line 3, which will require a single $O(n)$ pass to sum all relevant quantities via Welford’s online algorithm. ∎
Corollary 4.3.
Suppose that the Welford parameters of $D$ are computed in advance. Then $S_{1}\left(D,\mathcal{F}\right)$ can be computed in $O(1)$ time.
Proof.
We repeat our proof for Theorem 4.2, albeit excluding line 3, resulting in a net $O(1)$ runtime. ∎
Algorithm 1 Calculate maximum change in PCC by adding a new data point within a convex boundary
1:procedure ComputePrimarySensitivity($D,\mathcal{F}$)
2:     $S_{1}\leftarrow 0$
3:     $n,\bar{x}_{n},\bar{y}_{n},s_{x,n},s_{y,n},s_{xy,n}\leftarrow$ Welford parameters of $D$
4:     $\hat{\beta}_{x,n},\hat{\beta}_{y,n}\leftarrow\frac{s_{xy,n}}{s_{x,n}},\frac{s_{xy,n}}{s_{y,n}}$
5:     $\hat{\alpha}_{x,n},\hat{\alpha}_{y,n}\leftarrow\bar{y}_{n}-\hat{\beta}_{x,n},\bar{x}_{n}-\hat{\beta}_{y,n}\bar{y}_{n}$
6:     $r_{n}\leftarrow\frac{s_{xy,n}}{s_{x,n}s_{y,n}}$
7:     $\mathcal{C}\leftarrow\left\{(x,y)\mid(x=\ell_{x}\text{ or }x=u_{x})\text{ and }(y=\ell_{y}\text{ or }y=u_{y})\right\}$
8:     $\mathcal{I}_{n}\leftarrow\left\{\left(\hat{\beta}_{x,n}^{-1}\left(y-\hat{\alpha}_{x,n}\right),y\right)\mid y=\ell_{y}\text{ and }y=u_{y}\right\}$
9:     $\mathcal{I}_{n}\leftarrow\mathcal{I}_{n}\cup\left\{\left(x,\hat{\beta}_{y,n}^{-1}\left(x-\hat{\alpha}_{y,n}\right)\right)\mid x=\ell_{x}\text{ and }x=u_{x}\right\}$
10:     $\mathcal{I}_{n}\leftarrow\left\{(x,y)\in\mathcal{I}_{n}:\ell_{x}\leq x\leq u_{x}\text{ and }\ell_{y}\leq y\leq u_{y}\right\}$
11:     for each $(x_{n+1},y_{n+1})\in\mathcal{C}\cup\mathcal{I}_{n}$ do
12:         $\Delta x,\Delta y\leftarrow x_{n+1}-\bar{x}_{n},y_{n+1}-\bar{y}_{n}$
13:         $r_{n+1}\leftarrow\frac{(n+1)s_{x,n}s_{y,n}+\Delta x\Delta y}{\sqrt{\left((n+1)s_{x,n}^{2}+\Delta x^{2}\right)\left((n+1)s_{y,n}^{2}+\Delta y^{2}\right)}}$
14:         $\Delta r\leftarrow\left|r_{n}-r_{n+1}\right|$
15:         $S_{1}\leftarrow\max(S_{1},\Delta r)$
16:     end for
17:     return $S_{1}$
18:end procedure
5 Experiments
5.1 Synthetic data
Figure 3: Local sensitivities of the PCC for various distributions as computed by both ComputePrimarySensitivity and brute-force grid search. Our method consistently predicts a value at least as high as brute-force.
In order to empirically test the correctness of our algorithm, we generated synthetic datasets of varying sizes and compared the results of ComputePrimarySensitivity to brute-force grid search for the four following distributions. First, we sampled each coordinate from an independent uniform distributions along the interval $[-10,10]$. Second, we sampled from bivariate Gaussians with centroids at $(0,0)$ and covariance matrices given by $A^{\intercal}A$, where $A\in\left[0,1\right]^{2\times 2}$ is randomly generated from a uniform distribution at each iteration (multiplication by its own transpose was intended to enforce symmetry and positive definiteness). Third, we sampled from Dirichlet distributions with parameters generated by uniformly sampling $\boldsymbol{\alpha}\in\left[0,10\right]^{3}$. Fourth, we generated bivariate Gaussians and substituted 10% of each sample with outliers generated from a uniform distribution along the interval $\left[-30,30\right]$.
Each sample was generated so as to be of sizes $N\in\left\{10,50,100\right\}$, and for each distribution and $N$, we ran 100 separate experiments. The feasible region was simply set as the bounding box for each dataset, meaning that its size varied dynamically with each experiment. For grid search, we sampled 10 equidistant points along each axis of the feasible region for a total of 100 candidate points. Larger datasets and finer grids were not employed due to hardware limitations. All code experiments were conducted in Numpy [11].
Results (Figure 3) indicate that ComputePrimarySensitivity predicted sensitivities at least as high as those obtained by grid search at every iteration of our experiments. In 92.5% of the synthetic datasets, the sensitivities were nearly identical (i.e., within a relative tolerance of $1\times 10^{-5}$ to account for floating point error), with the sensitivities computed by ComputePrimarySensitivity slightly greater in each of the remaining 7.5% of cases. We conclude that our method correctly predicts the true sensitivity for every dataset, with the occasionally somewhat lower estimates returned by brute-force search simply reflecting the coarseness of the grid employed.
5.2 Real-world example: emerging markets
Among the many use cases for efficiently computing the sensitivity of the PCC, stock prices offer an excellent application due to both the inherently dynamic nature of financial data and the need for robust measures of collinearity. In the following example, we analyze the relationships between Alibaba (BABA) and the emerging markets (EEM) exchange-traded fund (ETF). We consider the opening price and trading volume for the instruments as sourced from Kaggle, restricting our analysis to the first 365 days available [4]. For price and volume, we compute the PCC between BABA and EEM and restrict our feasible region to be between 0 and twice the current maximum of the two stocks along their respective axes. In other words, the following experiment simulates an analysis in which we compute the extent to which the PCC for these variables can change by the end trading hours the following day, assuming that values will not fall below 0 or exceed double their previous maxima observed thus far. Once again, we compare the output of our method to that of grid search across 100 total points. Further, we applied a Fisher $z$-transformation to compute the lower and upper bounds of a 95% confidence interval [1].
Table 2: PCC, Confidence Intervals, and Predicted and Brute-Force Sensitivities
Metric PCC Lower Bound Upper Bound Predicted $S_{1}$ Brute-Force $S_{1}$
Opening Values 0.68734 0.62908 0.73793 0.37349 0.37349
Volume 0.13962 0.03750 0.23885 0.45219 0.45219
Results (Table 2) again indicate negligible differences between the sensitivities computed by brute-force and ComputePrimarySensitivity, indicating the validity of our algorithm on empirical datasets. Interestingly, the primary sensitivity is far greater than the size of the confidence intervals for both PCCs, underscoring the advantages of this metric over traditional hypothesis-testing for online applications in which outliers may be encountered.
6 Discussion
Practical applications of our approach are vast, extending to many fields in which dynamic datasets require reliable collinearity measures. In bioinformatics, the strength of associations between genetic markers and phenotypes in genome-wide association studies (GWAS) may be subject to change if data are mined from expanding biobanks [28, 30, 18]. In climatology, associations between temperature and various environmental outcomes like ground water availability are famously subject to reevaluation in light of ongoing studies [13, 29, 21]. In behavioral science and other experimental disciplines, similar analyses may be employed to halt experiments prematurely if the strength of association between variables is determined to be sufficiently robust to new subject data [3, 20, 25]. In the technology sector, dynamic environments like streaming platforms draw from real-time analytics to understand potentially unstable relationships between data streams, user preferences, internet traffic, etc. [22, 16] In differential privacy, the local sensitivity of metrics like the PCC is key for methods like Propose-Test-Release (PTR) [8].
As to the theoretical doors opened by our work, future directions are also numerous. First, in many optimization tasks like linear programming, feasible regions are often presumed to be generally polygonal (or, rather, polyhedral for higher-dimensional tasks) rather than simply rectangular [5]. We speculate that similar propositions like those proven above might be straightforwardly adapted for arbitrary piecewise linear boundaries, with basic feasible solutions playing roles similar to corner points in the case of a rectangular domain.
We conjecture that $k$-ary sensitivity can be efficiently computed as well by extending our theory further. In analyses not explicated in this paper, we were able to prove that the worst-case sequence of $k$ points to add to a bounded dataset must conform to certain structures, e.g., that each point in the sequence must also be a corner or intersection point. These patterns may offer an opportunity to compute the worst-case sequence via a combinatorial optimization approach.
Finally, our method of leveraging Welford’s identities may be extended to other linear metrics such as covariance and the parameters of the best linear unbiased estimators (BLUE) for simple linear regression [1]. For higher-dimensional datasets, similar analyses might be even extended to covariance matrices or principal components [1].
7 Conclusion
In this work, we develop both a rigorous theoretical framework for analyzing changes to the Pearson correlation coefficient (PCC) [23] induced by the addition of new data points and an algorithm for computing these changes in practice. Our key contributions consist of formulating a recurrence relation for the PCC via identities vital to Welford’s online algorithm [31], deriving closed-form solutions to this relation, and finally demonstrating that these solutions can be straightforwardly computed in linear or even constant time, depending on whether key parameters are precomputed. In addition to formal proofs of correctness, we validate our method on synthetic and experimental data to empirically confirm its efficacy.
Our code is made available at https://github.com/marc-harary/sensitivity for both practical usage and ongoing algorithmic development in this area.